OpenStudy (anonymous):

How do you know that tan(1/2) = pi/4 and tan(1/4) = 1.. etc. etc?

OpenStudy (lgbasallote):

hmm tan (1/2) = pi/4?

OpenStudy (anonymous):

my bad, tan(1) = pi/4. are there any tricks to remembering them? and I mean arctan!

OpenStudy (lgbasallote):

ahh are you familiar with the 45-45-90 triangle?

OpenStudy (lgbasallote):

 so as you can see from my triangle... \[\tan (\frac \pi 4 ) = 1\] therefore.. \[\tan^{-1} (1) = \frac \pi 4\] does that help?